OpenStudy (armyrangers):

If the sin 75° is 0.966, then the cos _____° is 0.966.

OpenStudy (armyrangers):

@adrynicoleb

OpenStudy (anonymous):

\[\cos ^{-1}(\sin (75 {}^{\circ}))=\frac{\pi }{12} \]\[\cos \left(\frac{\pi }{12}\right)=0.965926 \]

OpenStudy (armyrangers):

I take it as 12is the answer

OpenStudy (anonymous):

The answer is Pi/12 or 15 degrees.

OpenStudy (armyrangers):

Ok thanks

OpenStudy (anonymous):

You are welcome and thank you for the medal.